V1 = 0.55L, P1 = 740mmHg, V2 = 0.80 L, P2 =


1
Expert's answer
2021-04-27T07:48:14-0400

Given:

"V_1 = 0.55L, P_1 = 740mmHg, V_2 = 0.80 L, P_2=?"


We know that in a system moles are constant then the ideal gas equation seem as

"\dfrac{P_1}{V_1}=\dfrac{P_2}{V_2}\\"


"\Rightarrow P_2= \dfrac{P_1\times V_1}{V_2}\\"


"\Rightarrow P_2=\dfrac{740\times 0.55}{0.80}\\"


"\Rightarrow P_2=508.75mmHg\\"